An undiscovered planet, many lightyears from Earth, has one moon in a periodic orbit. This moon takes 1710 * 10^3 seconds (about 2.0 * 10^1 days) on average to complete one nearly circular revolution around the unnamed planet. If the distance from the center of the moon to the surface of the planet is 295.0 * 10^6 m and the planet has a radius of 3.20 ^ 10^6 m, calculate the moon\'s radial acceleration.
Answer #1

Distance between centres of mass, R = 2.95*10^6m + 3.20*10^6m

R = 6.15^6 m


Periodic time, T = 20d x 24h x 3600s

T = 1.728^6 s


Angular velocity, ? = 2? / T rad/s

=2?/1.72^6s

? = 3.651^-6 rad/s


Radial acceleration, a = R?